Radioactive nuclei has10 ^8 nuclei, it`s half life is 30s the number of nuclei left after 15s is

Grade:11

1 Answers

Vikas TU
6 years ago
 
Dear Student,
0.693/t1/2 = k
t1/2= half life 
k=disintegration or rot consistent 
=>0.693/30 = k
=>k=0.0231 
k=2.303log(a/a-x)/t 
a=initial sum 
a-x= sum left after deterioration 
t=time of rot 
=>0.0231=2.303log(108/108-x)/15 
=>108/108-x=1.41 
=>Amount left=108-x=7,09,21,986 nuclei.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)